# ABBA-Adapters: Efficient and Expressive Fine-Tuning of Foundation Models

### v3

@@ 1 Introduction
Large Language Models (LLMs) have become the backbone of modern NLP systems (1, 2, 3, 4, 5, 6, 7), demonstrating strong generalization across a wide range of tasks (8, 9). However, adapting these models to new tasks typically requires full fine-tuning (FT), which is computationally and memory intensive. Parameter-Efficient Fine-Tuning (PEFT) methods address this challenge by introducing a small number of trainable parameters while keeping the majority of model weights frozen (10, 11, 12). Among PEFT approaches, Low-Rank Adaptation (LoRA) (10) is the most widely adopted due to its simplicity and effectiveness. It models the weight update $\Delta W$ as the product of two low-rank matrices, providing a compact parameterization. However, this formulation inherently constrains updates to a low-dimensional subspace, limiting expressivity. Several extensions attempt to overcome this limitation: LoRA-XS (12) restricts updates to a frozen, SVD-derived subspace and training only a small higher-rank intermediary matrix while DoRA (11) modifies only the directional component of the weights via low-rank updates. Despite these architectural variations, all retain LoRA’s core constraint: updates remain strictly low-rank and limited in expressivity.
@@ 2.1 Preliminaries
LoRA (10). LoRA mitigates this by modeling the update as a low-rank decomposition: $\Delta W=sBA$, where $B\in\mathbb{R}^{m\times r}$, $A\in\mathbb{R}^{r\times n}$, and $s$ is a scaling factor. This reduces the number of trainable parameters to $r(m+n)$, with $r\ll\min(m,n)$. LoRA can represent any update of rank at most $r$, but cannot express higher-rank updates. Moreover, the projected gradient onto the weight space is also low-rank. While effective for simpler tasks, this limitation becomes significant in settings requiring high-rank updates or gradients (14, 15).
@@ 2.3 ABBA-Adapters: Improved Expressivity and High Rank, Yet Efficient
By the Eckart–Young–Mirsky (EYM) theorem (16, 17), the truncated SVD yields the optimal rank-$r_{1}$ approximation of $W_{0}$. This hybrid initialization anchors the update close to a meaningful low-rank subspace, while enabling the second adapter pair to explore task-specific directions during training. We validate the effectiveness of this strategy empirically in Section 4.1.
@@
Proof.
See Appendix D.1. ∎
To address this, we use Theorem 1 to rewrite ABBA in a LoRA-like form: let $B_{\text{kr}}=B_{1}\odot_{r}B_{2}$ and $A_{\text{kr}}=(A_{1}^{\top}\odot_{r}A_{2}^{\top})^{\top}$. The update becomes $\Delta Wx=B_{\text{kr}}(A_{\text{kr}}x)$, avoiding any full-rank construction. This enables ABBA to match LoRA’s compute and memory efficiency, while offering significantly higher expressivity, and remain more efficient than variants like HiRA, as shown in Section 4.4.
@@ 2.4 Expressivity of ABBA
and is lower-bounded by the classical EYM theorem (16, 17), which states that the optimal rank-$r$ approximation is given by the truncated SVD. Since a LoRA adapter of rank $r$ can only represent updates with rank at most $r$, EYM provides a theoretical minimum for $\mathcal{E}(r)$. LoRA achieves this bound exactly when the learned adapters align with the top singular components of $M$; otherwise, practical considerations such as suboptimal initialization may lead to a performance gap.
@@ 2.5 Stability of the ABBA Update
Definition 1 (Rank Stability of ABBA Adapters (21, 22)).
@@
Building on Definition 1, we establish in Theorem 2 that the ABBA update is indeed rank-stable.
@@ 3 Experiments
We evaluate ABBA on a range of models, specifically Llama-3.2 1B (23), Llama-3.2 3B (23), Mistral-7B (24), and Gemma-2 9B (25), to test its effectiveness across diverse scales and architectures. We run experiments on multiple benchmarks to capture varied trends. Appendix H details our training configurations, and Appendix I lists dataset specifics. For fair comparison with LoRA and other PEFT methods, we match the number of trainable parameters in ABBA by setting $r_{1}=r_{2}=r/2$.
@@ 4.5 Extending ABBA via Adapter Chains
A natural extension of ABBA is decomposing the update into a composition of $k$ multiple adapter pairs: $\Delta W=B_{1}A_{1}\odot B_{2}A_{2}\dots\odot B_{k}A_{k},$ where each adapter pair has rank $r/k$, with $r$ denoting the total rank budget. This factorized form increases the expressive capacity and the effective rank of the update. However, this introduces additional optimization challenges and may reduce training stability. We evaluate a variant of ABBA using a chain of four adapter pairs and compare it to the standard two-pair setup. As shown in Table 6, the chained version performs slightly worse, likely due to suboptimal scaling or training instability. We leave further investigation of multi-stage compositions to future work.
@@ Appendix A Related Work
PEFT methods adapt large pretrained models to downstream tasks by training a small number of additional parameters while keeping the base model frozen. Among these, LoRA (10) is widely adopted for its simplicity and effectiveness, modeling the update $\Delta W$ as a low-rank product $BA$, thereby reducing trainable parameters significantly. Several extensions enhance LoRA along different axes. QLoRA (38) and QA-LoRA (39) combines quantization with LoRA to reduce memory footprint; AdaLoRA (40) allocates rank budgets dynamically across layers. LoRA-XS (12) inserts a small trainable core between frozen adapters to improve compression, while LoRA-Pro (14) and LoRA-SB (15) optimize adapters to better approximate full fine-tuning gradients. Other variants modify structure or initialization. VeRA (41) reuses frozen adapters across layers with task-specific scaling vectors; DoRA (11) applies low-rank updates only to the direction of pretrained weights. PiSSA (26) initializes adapters using top singular vectors, while rsLoRA (21) proposes scale-aware initialization to improve stability. Despite all these differences, these methods share a core principle: they represent $\Delta W$ using a low-rank structure, enabling adaptation under tight compute and memory constraints. LoRA-based methods have also been applied in other domains, such as federated fine-tuning (42, 20, 43, 44).
@@
While low-rank PEFT methods offer strong efficiency–performance tradeoffs, they can underperform in tasks requiring high-rank updates. This limitation has driven recent efforts to move beyond purely low-rank parameterizations. HiRA (13) achieves high-rank updates by taking the Hadamard (elementwise) product of the pretrained matrix $W_{0}$ with a low-rank adapter $BA$, leveraging the fact that such a product can increase effective rank without increasing parameter count. MoRA (45) instead learns a full-rank update through input compression and activation decompression, while KronA (46) uses Kronecker products between adapters to boost representational capacity. ReLoRA (47) uses multiple low-rank updates to approximate a final higher-rank update. Other approaches explore elementwise structure for different purposes: FLoRA (42) modulates intermediate activations per task using Hadamard products, and PACE (48) injects multiplicative noise during adaptation. These trends reflect a broader shift toward structured, high-rank updates for improved expressivity.
@@ Appendix B Why Does No Closed‑Form Reconstruction Solution Exist for ABBA?
to decouple along singular directions, as guaranteed by the Eckart–Young–Mirsky theorem (16, 17).
@@ Appendix D Proofs
D.1 Proof of Theorem 1: Khatri–Rao Factorization
Theorem (Khatri–Rao Factorization (18)).
Let $B_{1}A_{1},B_{2}A_{2}\in\mathbb{R}^{m\times n}$ . Then, $(B_{1}A_{1})\odot(B_{2}A_{2})=\underbrace{(B_{1}\odot_{r}B_{2})}_{m\times r_{1}r_{2}}\underbrace{(A_{1}^{\top}\odot_{r}A_{2}^{\top})^{\top}}_{r_{1}r_{2}\times n},$ where $\odot_{r}$ denotes the row-wise Khatri–Rao product.
Proof.
We prove the equality by comparing the $(i,j)$-th elements of the left-hand side (LHS) and the right-hand side (RHS). Starting with the LHS, we have:
$[(B_{1}A_{1})\odot(B_{2}A_{2})]_{ij}$ $=(B_{1}A_{1})_{ij}\cdot(B_{2}A_{2})_{ij}$
$=\left(\sum_{k=1}^{r_{1}}B_{1,ik}A_{1,kj}\right)\cdot\left(\sum_{l=1}^{r_{2}}B_{2,il}A_{2,lj}\right)$
$=\sum_{k=1}^{r_{1}}\sum_{l=1}^{r_{2}}B_{1,ik}A_{1,kj}B_{2,il}A_{2,lj}.$
Next, we analyze the RHS. Observe that $(B_{1}\odot_{r}B_{2})$ is an $m\times(r_{1}r_{2})$ matrix, and $(A_{1}^{\top}\odot_{r}A_{2}^{\top})$ is an $n\times(r_{1}r_{2})$ matrix. The matrix product on the RHS is:
$\left[(B_{1}\odot_{r}B_{2})(A_{1}^{\top}\odot_{r}A_{2}^{\top})^{\top}\right]_{ij}$ $=\sum_{f=1}^{r_{1}r_{2}}\left[B_{1}\odot_{r}B_{2}\right]_{if}\cdot\left[A_{1}^{\top}\odot_{r}A_{2}^{\top}\right]_{jf}.$
To evaluate this sum, we define indices $k:=\left\lfloor\frac{f-1}{r_{2}}\right\rfloor+1$ and $l:=(f-1)\bmod r_{2}+1$, so that $f$ ranges over all $(k,l)$ pairs with $k\in[1,r_{1}]$ and $l\in[1,r_{2}]$. Then:
$\left[B_{1}\odot_{r}B_{2}\right]_{if}$ $=B_{1,ik}B_{2,il},$
$\left[A_{1}^{\top}\odot_{r}A_{2}^{\top}\right]_{jf}$ $=A_{1,kj}A_{2,lj}.$
Substituting back into the sum:
$\sum_{f=1}^{r_{1}r_{2}}\left[B_{1}\odot_{r}B_{2}\right]_{if}\cdot\left[A_{1}^{\top}\odot_{r}A_{2}^{\top}\right]_{jf}$ $=\sum_{k=1}^{r_{1}}\sum_{l=1}^{r_{2}}B_{1,ik}B_{2,il}A_{1,kj}A_{2,lj}.$
This expression matches exactly with the earlier expansion of the LHS. Hence, the two sides are equal, completing the proof. ∎
Given two matrices $U,V\in\mathbb{R}^{m\times n}$, we define the row-wise Hadamard product, denoted by $U\odot_{r}V$. Let $U_{i}$ and $V_{i}$ denote the $i$-th rows of $U$ and $V$, respectively. The row-wise Hadamard product is computed as:
$[U\odot_{r}V]_{i}:=[U_{i1}\cdot V_{i},\,U_{i2}\cdot V_{i},\,\dots,\,U_{in}\cdot V_{i}],$
where each element of row $U_{i}$ scales the entire corresponding row $V_{i}$.
Consider the following example:
$U=\begin{bmatrix}1&2\\ 3&4\end{bmatrix},\quad V=\begin{bmatrix}3&4\\ 5&6\end{bmatrix}$.
Applying the row-wise Hadamard product $U\odot_{r}V$, we compute:
$[U\odot_{r}V]_{1}$ $:=[1\cdot V_{1},\;2\cdot V_{1}]=\begin{bmatrix}3&4&6&8\end{bmatrix}$,
$[U\odot_{r}V]_{2}$ $:=[3\cdot V_{2},\;4\cdot V_{2}]=\begin{bmatrix}15&18&20&24\end{bmatrix}$.
Combining both rows, we obtain the final result:
$U\odot_{r}V$ $=\begin{bmatrix}3&4&6&8\\ 15&18&20&24\end{bmatrix}$.
D.2 Proof of Theorem 2: Rank-Stability of ABBA
@@ Appendix G Training Time
Appendix H Experimental Details
@@
Appendix I Dataset Details
@@
Appendix J Use of Large Language Models

### v4

@@ 1 Introduction
Large Language Models (LLMs) have become the backbone of modern NLP systems (34, 22, 1, 45, 42, 51, 43), demonstrating strong generalization across a wide range of tasks (4, 14). However, adapting these models to new tasks typically requires full fine-tuning (FT), which is computationally and memory intensive. Parameter-Efficient Fine-Tuning (PEFT) methods address this challenge by introducing a small number of trainable parameters while keeping the majority of model weights frozen (17, 25, 2). Among PEFT approaches, Low-Rank Adaptation (LoRA) (17) is the most widely adopted due to its simplicity and effectiveness. It models the weight update $\Delta W$ as the product of two low-rank matrices, providing a compact parameterization. However, this formulation inherently constrains updates to a low-dimensional subspace, limiting expressivity. Several extensions attempt to overcome this limitation: LoRA-XS (2) restricts updates to a frozen, SVD-derived subspace and training only a small higher-rank intermediary matrix while DoRA (25) modifies only the directional component of the weights via low-rank updates. Despite these architectural variations, all retain LoRA’s core constraint: updates remain strictly low-rank and limited in expressivity.
@@ 2.1 Preliminaries
LoRA (17). LoRA mitigates this by modeling the update as a low-rank decomposition: $\Delta W=sBA$, where $B\in\mathbb{R}^{m\times r}$, $A\in\mathbb{R}^{r\times n}$, and $s$ is a scaling factor. This reduces the number of trainable parameters to $r(m+n)$, with $r\ll\min(m,n)$. LoRA can represent any update of rank at most $r$, but cannot express higher-rank updates. Moreover, the projected gradient onto the weight space is also low-rank. While effective for simpler tasks, this limitation becomes significant in settings requiring high-rank updates or gradients (47, 33).
@@ 2.3 ABBA-Adapters: Improved Expressivity and High Rank, Yet Efficient
By the Eckart–Young–Mirsky (EYM) theorem (12, 29), the truncated SVD yields the optimal rank-$r_{1}$ approximation of $W_{0}$. This hybrid initialization anchors the update close to a meaningful low-rank subspace, while enabling the second adapter pair to explore task-specific directions during training. We validate the effectiveness of this strategy empirically in Section 4.1.
@@
To address this, we use Theorem 1 to rewrite ABBA in a LoRA-like form: let $B_{\text{kr}}=B_{1}\odot_{r}B_{2}$ and $A_{\text{kr}}=(A_{1}^{\top}\odot_{r}A_{2}^{\top})^{\top}$. The update becomes $\Delta Wx=B_{\text{kr}}(A_{\text{kr}}x)$, avoiding any full-rank construction.
This enables ABBA to match LoRA’s compute and memory efficiency, while offering significantly higher expressivity, and remain more efficient than variants like HiRA, as shown in Section 4.4.
@@ 2.4 Expressivity of ABBA
and is lower-bounded by the classical EYM theorem (12, 29), which states that the optimal rank-$r$ approximation is given by the truncated SVD. Since a LoRA adapter of rank $r$ can only represent updates with rank at most $r$, EYM provides a theoretical minimum for $\mathcal{E}(r)$. LoRA achieves this bound exactly when the learned adapters align with the top singular components of $M$; otherwise, practical considerations such as suboptimal initialization may lead to a performance gap.
@@ 2.5 Stability of the ABBA Update
Definition 1 (Rank Stability of ABBA Adapters (21, 46)).
@@
Building on Definition 1, we establish in Theorem 2 that the ABBA update is indeed rank-stable. We empirically demonstrate that our scaling yields improved stability, as evidenced by the gradient norm experiments shown in Figure 5 in Appendix I.
@@ 3 Experiments
We evaluate ABBA on a range of models, specifically Llama-3.2 1B (11), Llama-3.2 3B (11), Mistral-7B (38), and Gemma-2 9B (44), to test its effectiveness across diverse scales and architectures. We run experiments on multiple benchmarks to capture varied trends. Appendix L details our training configurations, and Appendix M lists dataset specifics. For fair comparison with LoRA and other PEFT methods, we match the number of trainable parameters in ABBA by setting $r_{1}=r_{2}=r/2$.
@@ 4.4 All About Efficiency
4.5 Additional Results on Arithmetic Tasks
We report results for fine-tuning Mistral-7B and Gemma-2 9B using a larger 40K subset of MetaMathQA, compared to the 20K subset used in the main experiments.
As shown in Table 6, ABBA continues to outperform all other PEFT baselines under matched parameter budgets. However, we observe that the performance gap between full fine-tuning and ABBA narrows when using the larger subset. This suggests that in higher data regimes, full fine-tuning may be a more suitable choice relative to other PEFT methods, provided the additional memory and compute requirements for full fine-tuning can be met.
Table 6: Comparison of multiple FT methods on Mistral-7B and Gemma-2 9B across arithmetic reasoning benchmarks (when finetuned on a 40K subset of MetaMathQA). Best results among PEFT methods are in bold.
Method Mistral-7B Gemma-2 9B
Full FT $7.24$ B $66.28$ $18.34$ $9.24$ B $79.89$ $39.44$
LoRA $83.88$ M $62.89$ $16.45$ $108.04$ M $77.02$ $37.05$
DoRA $85.26$ M $63.24$ $16.67$ $109.88$ M $77.46$ $37.40$
HiRA $83.88$ M $64.54$ $17.88$ $108.04$ M $78.90$ $38.53$
ABBAr=16 $41.94$ M $66.36$ $18.51$ $54.02$ M $79.93$ $39.27$
ABBAr=32 $83.88$ M 67.04 18.76 $108.04$ M 80.31 39.92
@@ 4.6 Extending ABBA via Adapter Chains
A natural extension of ABBA is decomposing the update into a composition of $k$ multiple adapter pairs: $\Delta W=B_{1}A_{1}\odot B_{2}A_{2}\dots\odot B_{k}A_{k},$ where each adapter pair has rank $r/k$, with $r$ denoting the total rank budget. This factorized form increases the expressive capacity and the effective rank of the update. However, this introduces additional optimization challenges and may reduce training stability. We evaluate a variant of ABBA using a chain of four adapter pairs and compare it to the standard two-pair setup. As shown in Table 7, the chained version performs slightly worse, likely due to suboptimal scaling or training instability. We show gradient norm plots of these variations in Figure 6 in Appendix J. We leave further investigation of multi-stage compositions to future work.
@@ 5 Conclusion
Reproducibility Statement
We have made considerable efforts to ensure that our work is fully reproducible. Our code is open-source and available at: https://github.com/CERT-Lab/abba. Detailed descriptions of our experimental settings can be found in Section 3 and Appendix L. All datasets in this work are widely used, publicly available benchmarks (see Appendix M for details).
@@ Appendix A Related Work
PEFT methods adapt large pretrained models to downstream tasks by training a small number of additional parameters while keeping the base model frozen. Among these, LoRA (17) is widely adopted for its simplicity and effectiveness, modeling the update $\Delta W$ as a low-rank product $BA$, thereby reducing trainable parameters significantly. Several extensions enhance LoRA along different axes. QLoRA (10) and QA-LoRA (50) combines quantization with LoRA to reduce memory footprint; AdaLoRA (54) allocates rank budgets dynamically across layers. LoRA-XS (2) inserts a small trainable core between frozen adapters to improve compression, while LoRA-Pro (47) and LoRA-SB (33) optimize adapters to better approximate full fine-tuning gradients. Other variants modify structure or initialization. VeRA (23) reuses frozen adapters across layers with task-specific scaling vectors; DoRA (25) applies low-rank updates only to the direction of pretrained weights. PiSSA (27) initializes adapters using top singular vectors, while rsLoRA (21) proposes scale-aware initialization to improve stability. LoRA-One (55) initializes adapters using this singular subspaces of the full fine-tuning gradient and incorporates preconditioning to improve convergence. WeGeFT (37) learns a small low-rank transformation that maps pretrained weights to their fine-tuned updates, either shared across layers or learned per layer. Despite all these differences, these methods share a core principle: they represent $\Delta W$ using a low-rank structure, enabling adaptation under tight compute and memory constraints. LoRA-based methods have also been applied in other domains, such as federated fine-tuning (48, 30, 40, 39).
@@
While low-rank PEFT methods offer strong efficiency–performance tradeoffs, they can underperform in tasks requiring high-rank updates. This limitation has driven recent efforts to move beyond purely low-rank parameterizations. HiRA (19) achieves high-rank updates by taking the Hadamard (elementwise) product of the pretrained matrix $W_{0}$ with a low-rank adapter $BA$, leveraging the fact that such a product can increase effective rank without increasing parameter count. MoRA (20) instead learns a full-rank update through input compression and activation decompression, while KronA (13) uses Kronecker products between adapters to boost representational capacity. ReLoRA (24) uses multiple low-rank updates to approximate a final higher-rank update. Other approaches explore elementwise structure for different purposes: FLoRA (wang2024florafederatedfinetuninglarge) modulates intermediate activations per task using Hadamard products, and PACE (31) injects multiplicative noise during adaptation. These trends reflect a broader shift toward structured, high-rank updates for improved expressivity.
@@ Appendix B Why Does No Closed‑Form Reconstruction Solution Exist for ABBA?
to decouple along singular directions, as guaranteed by the Eckart–Young–Mirsky theorem (12, 29).
@@ Appendix C Why Does LoRA-Style Initialization of Both Adapter Pairs Fail?
Appendix D Proof of Theorem 2: Rank-Stability of ABBA
@@ Appendix G Training Time
Appendix H Performance on Coding Tasks
We also evaluate ABBA against other PEFT variants on code generation tasks. We fine-tuned Llama-3.2-1B on a 100K subset of the CodeFeedback dataset (56) and assessed performance using Pass@1 on HumanEval (5). As shown in Table 11, ABBA again performs very strongly, outperforming all other PEFT variants and even full fine-tuning (all PEFT methods are run at rank 32). These results demonstrate that the advantages of our method generalize to more complex tasks such as code generation.
Table 11: Comparison of multiple FT methods across the coding benchmark - HumanEval. Best results among PEFT methods are in bold.
Method # Params HumanEval ($\uparrow$)
Full FT $1.24$ B $23.17$
LoRA $22.54$ M $20.73$
HiRA $22.54$ M $21.95$
ABBA $22.54$ M $\mathbf{25.61}$
Appendix I Rank-Stability of ABBA: Gradient Norms
In Theorem 2, we had established the optimal scaling for ABBA,
$s_{\text{ABBA}}\in\Theta\!\left(\tfrac{1}{\sqrt{r_{1}r_{2}}}\right).$
We now verify this result empirically by comparing several scaling strategies:
$s_{\text{ABBA}}\in\Theta(1),\quad\Theta\!\left(\tfrac{1}{(r_{1}r_{2})^{1/6}}\right),\quad\Theta\!\left(\tfrac{1}{(r_{1}r_{2})^{1/4}}\right),\quad\Theta\!\left(\tfrac{1}{(r_{1}r_{2})^{1/2}}\right).$
As shown in Figure 5, the scaling $s_{\text{ABBA}}\in\Theta\!\left(\tfrac{1}{\sqrt{r_{1}r_{2}}}\right)$ produces the most stable gradient norms, consistent with our theoretical findings.
Figure 5: Gradient norms in Mistral-7B comparing various scaling strategies.
Appendix J Stability in Chained ABBA: Gradient Norms
We present gradient norms for the chained ABBA setup (with four adapter pairs) compared to the conventional ABBA configuration (with two pairs) in Figure 6. As hypothesized in Section 4, the chained variant exhibits significantly greater gradient-norm instability than the standard ABBA model.
Figure 6: Gradient norms in Mistral-7B comparing chained ABBA (4 pairs) and normal ABBA
Appendix K Comparison with High-Rank Baselines
We evaluate ABBA against additional higher-rank PEFT variants, namely Flora (15), MoRA (jiang2024morahighrankupdatingparameterefficient), and KronA (13), on code generation tasks. We fine-tuned Llama-3.1 8B on a 100K subset of the CodeFeedback dataset (56) and assessed performance using Pass@1 on HumanEval (5). As shown in Table 12, ABBA again performs very strongly, outperforming all other higher-rank PEFT variants.
Table 12: Comparison of multiple PEFT methods across the coding benchmark - HumanEval. Best results among PEFT methods are in bold.
Method HumanEval ($\uparrow$)
Full FT $48.54$
LoRA $45.73$
KronA $45.12$
Flora $45.56$
MoRA $47.28$
HiRA $48.17$
ABBA $\mathbf{49.39}$
Appendix L Experimental Details
@@ Appendix I Rank-Stability of ABBA: Gradient Norms
Appendix M Dataset Details
@@
Appendix N Use of Large Language Models